# ContextualStory: Consistent Visual Storytelling with Spatially-Enhanced and Storyline Context

### v2

@@ Abstract
Visual storytelling involves generating a sequence of coherent frames from a textual storyline while maintaining consistency in characters and scenes. Existing autoregressive methods, which rely on previous frame-sentence pairs, struggle with high memory usage, slow generation speeds, and limited context integration. To address these issues, we propose ContextualStory, a novel framework designed to generate coherent story frames and extend frames for story continuation. ContextualStory utilizes Spatially-Enhanced Temporal Attention to capture spatial and temporal dependencies, handling significant character movements effectively. Additionally, we introduces a Storyline Contextualizer to enrich context in storyline embedding and a StoryFlow Adapter to measure scene changes between frames for guiding model. Extensive experiments on PororoSV and FlintstonesSV benchmarks demonstrate that ContextualStory significantly outperforms existing methods in both story visualization and story continuation.
@@ Introduction
Recent text-to-image (T2I) models, such as SD3 (Esser et al. 2024), excel at generating images from text but only produce individual images independently. Although text-to-video (T2V) models like SVD (Blattmann et al. 2023a) and Sora (Brooks et al. 2024) create temporally coherent videos from text, these typically feature simple scene or motion changes. In contrast, this paper focuses on visual storytelling, which comprises generating a sequence of coherent story frames from a textual storyline in story visualization and extending an initial frame from a textual storyline in story continuation. This task has significant potential for educational applications, such as crafting vivid, coherent comics for storybooks. The key challenge is aligning generated frames with sentences while ensuring temporal consistency in characters and scenes throughout the storyline. Providing sufficient context is essential due to the limited information in individual sentences.
@@
Figure 1: Story frames generated by our ContextualStory model on the PororoSV dataset. SETA and SC enhance character consistency and scene coherence, reducing issues like repeated characters. Compared to AR-LDM, ContextualStory achieves superior consistency across frames.
@@
Many diffusion-based visual storytelling methods use an autoregressive generative approach to capture temporal dependencies based on previous frame-sentence pairs to generate consistent story frames, such as AR-LDM (Pan et al. 2024) and Story-LDM (Rahman et al. 2023). However, these methods face four key limitations: 1) High memory usage due to storing all previous image-sentence pairs, making longer storyline difficult to handle; 2) Limited context in early image generation, which may impact image quality; 3) Slow generation speed due to the sequential nature of the process; 4) Inconsistency in generated images because they rely only on past pairs and neglect future context, missing the global story context. We address this by exploring how the model can 1) access sufficient frame context and 2) obtain adequate context from the storyline.
@@
To access sufficient frame context, we integrate temporal convolutions and Spatially-Enhanced Temporal Attention (SETA) into the UNet, combining them with the existing spatial modeling layer. By alternating between spatial and temporal modeling, the model can effectively capture spatial dependencies within individual frames and temporal dependencies across frames, allowing for comprehensive context from all frames. To obtain adequate context from the storyline, we propose the Storyline Contextualizer (SC), which further processes the text embeddings from the CLIP text encoder. This ensures that the context information from the storyline is propagated to each sentence, providing sufficient context throughout.
@@
The temporal attention is crucial for visual storytelling as it enables the propagation of context across story frames along the temporal dimension. However, vanilla temporal attention struggles to address the common issue of significant character movement across frames, as shown in Figure 3(a). To overcome this, we propose SETA that employs a local window mechanism across other frames. This allows the query to attend features within the local windows of other frames, effectively capturing characters with significant movement and enhancing the spatial dependency capture capability of temporal attention. By comparing the story frames generated by ContextualStory w/o SETA & SC and ContextualStory w/o SC in Figure 1, it is evident that SETA improves character consistency and reduces the occurrence of repeated characters, such as Pororo.
@@
The Storyline Contextualizer enhances the contextual information of storyline embeddings extracted from the CLIP text encoder, which initially contain only sentence-level semantics. By integrating and propagating context across all sentences, it generates a context-enriched storyline embedding. The Storyline Contextualizer is a transformer-based network that incorporates self-attention and temporal attention layers to capture both global and temporal dependencies. This enriched embedding is then used to guide the model through a temporally-aligned cross-attention mechanism, ensuring consistent story frames throughout the storyline. By comparing the story frames generated by ContextualStory w/o SC and those with SC in Figure 1, we observe that SC promotes the creation of more consistent and coherent scenes. Additionally, compared to the story frames generated by AR-LDM, our ContextualStory significantly enhances the consistency of both characters and scenes.
@@
Additionally, to leverage scene changes between story frames to guide the model, we proposed the StoryFlow Adapter to measure change between story frames. For story continuation, We simply add a convolution layer at the input end of the UNet block to match the size of the first frame latent with noise latent, and then concatenated them.
@@
Our contributions are as follows: (1) ContextualStory Framework: Our novel framework overcomes limitations of existing autoregressive methods, which face issues like high memory usage, limited context, slow generation speed, and image inconsistency. (2) Spatially-Enhanced Temporal Attention (SETA): We present the SETA into the UNet model, combining temporal convolutions with spatial modeling to capture both spatial and temporal dependencies, addressing challenges of significant character movement and improving consistency across frames. (3) Storyline Contextualizer (SC): This module is proposed processing text embeddings from the CLIP text encoder to enrich context across the entire storyline. It utilizes a transformer-based network to capture global and temporal dependencies, ensuring consistent and coherent story frames. (4) StoryFlow Adapter: We repurpose this tool to measure scene changes between story frames to guide the model in handling scene transitions more effectively. (5) Extensive experiments on the PororoSV and FlintstonesSV datasets demonstrate that our ContextualStory significantly outperforms previous SOTA in visual storytelling tasks.
@@
Visual storytelling. Early methods for story visualization primarily relied on GANs (Goodfellow et al. 2020). StoryGAN (Li et al. 2019) pioneers story visualization with a sequential conditional GAN framework, featuring a context encoder and dual discriminators for enhanced narrative and visual coherence. Many subsequent works (Song et al. 2020; Li, Torr, and Lukasiewicz 2022; Maharana, Hannan, and Bansal 2021; Maharana and Bansal 2021; Li 2022) improve StoryGAN, while others (Ahn et al. 2023; Chen et al. 2022) adopt Transformer-based methods to enhance character consistency. StoryDALL-E (Maharana, Hannan, and Bansal 2022) extends the story visualization to story continuation with a given initial frame and uses pre-trained DALL-E (Ramesh et al. 2021) to generate coherent images. Recently, diffusion models (DM) (Ho, Jain, and Abbeel 2020) have shown success in image generation. Some works (Pan et al. 2024; Rahman et al. 2023; Feng et al. 2023; Song et al. 2024; Liu et al. 2024; Shen and Elhoseiny 2023; Wang et al. 2024) propose an autoregressive diffusion framework based on previous captions and generated frames to enhance consistency. For example, Story-LDM (Rahman et al. 2023) incorporates a visual memory module to capture context of previous generated images. However, these autoregressive methods are memory-intensive and often fail to capture the global context of the storyline, leading to reduced overall consistency. RCDMs (Shen et al. 2024) is a two-stage model that first predicts the embedding of the unknown clip and then generates the corresponding images. StoryImager (Tao et al. 2024) is a unified framework capable of performing Story Visualization, Story Continuation, and Story Completion tasks. StoryGPT-V (Shen and Elhoseiny 2023) combines the image generation capability of LDM with the reasoning ability of LLM to ensure semantic consistency. TaleCrafter (Gong et al. 2023), Animate-A-Story (He et al. 2023), and AutoStory (Wang et al. 2023c) focus on designing system pipelines for story visualization, all employing LLM to generate storylines. In contrast, our ContextualStory addresses consistency by leveraging spatial-temporal attention to learn complex dependencies in images, departing from autoregressive methods.
@@
Text-to-image generation. Recently, significant progress (Rombach et al. 2022; Podell et al. 2023; Saharia et al. 2022; Ramesh et al. 2022) has been achieved in T2I generation, primarily due to advancements in DM (Ho, Jain, and Abbeel 2020). Another line of work (Liu et al. 2023; Dhariwal and Nichol 2021; Ho and Salimans 2022; Gal et al. 2022; Ruiz et al. 2023; Kumari et al. 2023) focuses on flexible and controllable image generation, including ControlNet (Zhang, Rao, and Agrawala 2023), Composer (Huang et al. 2023), IP-Adapter (Ye et al. 2023), and T2I-Adapter (Mou et al. 2024). ControlNet provides a general pipeline for conditioning on both text and image data. The Diffusion Transformer (Peebles and Xie 2023) showcases scalability by replacing UNet with a Transformer, and Pixart-$\alpha$ (Chen et al. 2023) further reduces training costs while achieving superior image quality. However, these methods focus on generating individual image aligned with text and struggle to produce multiple coherent and consistent images in a sequence.
@@
Diffusion models (DM) (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) are generative models that approximate data distributions by iteratively denoising a Gaussian distribution through a reverse process of a Markov Chain. Given a training sample $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ and add Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to the input in a forward process $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}:={\textstyle\prod_{s=1}^{t}}(1-\beta_{s})$ and $\beta_{1},\ldots,\beta_{T}$ is the variance schedule. The model is trained to approximate the backward process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ by minimizing the mean squared error (MSE) between predicted and target noise:
$L_{DM}:=\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|^{2}\right].$ (1)
Latent Diffusion Models (LDM) (Rombach et al. 2022) extend DM to high-dimensional data by compressing images into latent space. An encoder $\mathcal{E}$ maps the input $\mathbf{x}$ to a latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})$, and the forward and backward processes are applied on $\mathbf{z}$. The denoising network $\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ is trained by minimizing
$L_{LDM}:=\mathbb{E}_{\mathcal{E}(\mathbf{x}),\mathbf{c},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})\|^{2}\right],$ (2)
where $\mathbf{c}$ denotes conditional signals, such as storyline embeddings. The generated image $\mathbf{\hat{x}}$ is obtained by decoding the denoised latent $\mathbf{z}$ with pre-trained decoder $\mathcal{D}(\mathbf{z})$.
@@
Previous methods based on text-to-image (T2I) diffusion models typically use an autoregressive approach, generating each story frame sequentially rely on the preceding frames and captions. However, these methods often fail to capture sufficient storyline context, leading to poor consistency across the generated story frames. Moreover, the UNet of T2I models struggles to capture temporal dependencies between story frames, and vanilla temporal attention layers are ineffective in addressing large character movement across frames. To overcome these challenges, as illustrated in Figure 2, we introduce temporal convolution and Spatially-Enhanced Temporal Attention into the UNet. These additions help the model capture contextual information across both spatial and temporal dimensions, allowing it to better handle complex spatial and temporal dependencies. We also propose Storyline Contextualizer to ensure that the contextual information of the storyline propagates to each sentence. Additionally, to address the significant changes in characters and scenes between story frames, we introduce the StoryFlow Adapter to quantify these changes and guide the model in generating more coherent visual stories.
@@
Spatially-Enhanced Temporal Attention. In video diffusion models (Zhang et al. 2023), temporal attention layers are often employed to model temporal dependencies. However, story frames differ significantly from video frames, where characters and scenes tend to change considerably, whereas video frames, even keyframes, exhibit minimal changes and contain many redundant pixels. As illustrated in Figure 3, story frames often involve significant character movement, making it difficult for vanilla temporal attention to capture the same character across story frames.
To address this challenge, we propose Spatially-Enhanced Temporal Attention. Assuming the green block within the red-bordered area is the query, the query itself, along with the green blocks covered by the $k\times k$ local window at the same position in other frames (i.e., all the green blocks), form the key and value. Formally, given an hidden state $\mathcal{Z}_{t}=\{\mathbf{z^{1}_{t}},\dots,\mathbf{z^{N}_{t}}\}\in\mathbb{R}^{n\times c\times h\times w}$, where $n=N$ is the number of story frames, and $c$, $h$, $w$ represent the channel, height, and width dimensions of the hidden state, respectively. We first reshape it to $\mathcal{Z}_{t}^{\prime}\in\mathbb{R}^{hw\times n\times c}$, then extract the local window feature $\mathcal{Z}_{t}^{lw^{\prime}}\in\mathbb{R}^{hw\times n^{lw}\times c}$ at each spatial position, where $n^{lw}=(n-1)k^{2}+1$. Subsequently, we compute the query, key, and value and then perform the $\mathrm{Attention}(Q_{T},K_{T},V_{T})$ through Eq. (4).
@@
where $W_{T}^{Q}$, $W_{T}^{K}$, $W_{T}^{V}$ are learnable projection matrices, and $d$ is the feature dimensionality. To ensure that the features at the boundaries have a complete local window, we pad $\mathcal{Z}_{t}^{\prime}$ by replicating the boundary features. We utilized rotary positional embedding (RoPE) (Su et al. 2024) as the temporal positional embedding to enable the model to understand the temporal relationships between frames.
@@
Table 2: Quantitative comparison with SOTA methods of story visualization on PororoSV.
Storyline Contextualizer. We use the CLIP text encoder to independently extract text embeddings for each sentence in the storyline. These embeddings contain the semantic information of the corresponding sentences but lack the global contextual information of the storyline. Directly using these text embeddings to guide the model may result in inconsistent story frames. To address this challenge, we propose the Storyline Contextualizer, which propagates and integrates the contextual information from all text embeddings to create a context-enriched storyline embedding. As shown in Figure 2, the Storyline Contextualizer is a transformer-based network, each layer contains a self-attention layer, a temporal attention layer, and two Feed-Forward Network (FFN).
Given the storyline embedding $\mathcal{C}=\{\mathbf{c^{1}},\dots,\mathbf{c^{N}}\}\in\mathbb{R}^{n\times l\times c_{T}}$ from the CLIP text encoder, where $n=N$ is the number of story sentences, $l$ is the length of the feature sequence, and $c_{T}$ is the feature dimension. In Storyline Contextualizer, we first reshape the storyline embedding to $1\times nl\times c_{T}$, then perform self-attention. Then reshape the embedding to $l\times n\times c_{T}$ , and then perform temporal attention. To minimize any adverse effects from the additional modules, we zero-initialize the weights of the second FFN in the final layer and incorporate a residual connection, ensuring that the Storyline Contextualizer functions as an identity mapping at the beginning of training. The context-enriched storyline embedding $\mathcal{C}^{\prime}$ guides the model through a cross-attention layer. Unlike T2I and T2V models, which calculate cross-attention between image/frames and a single text, we adopt a temporally-aligned approach, computing cross-attention between each individual context-enriched text embedding $\mathbf{c^{i}}$ and hidden state $\mathbf{z^{i}_{t}}$ of the corresponding frame.
StoryFlow adapter. To leverage scene changes between story frames to guide the model, we propose the StoryFlow Adapter, inspired by (Jeong, Park, and Ye 2023; Wu et al. 2023; Qing et al. 2023). As shown in Figure 2, the storyflow is computed as the L2 norm of the differences between adjacent images, $\delta_{i}=\left\|\mathbf{z_{0}^{i}}-\mathbf{z_{0}^{i+1}}\right\|$, to quantify the image difference. Given $N$ ground truth frames, we calculate the storyflow $\Delta=\{\delta^{1},\dots,\delta^{N-1}\}$. We then encode the storyflow $\Delta$ into a $c$-dimensional embedding using sinusoidal embedding and a zero-initialized FFN. Through linear interpolation, we obtain the storyflow embedding $\Delta^{\prime}\in\mathbb{R}^{N\times c}$. Finally, we add storyflow embedding to timestep embedding and feed them into the spatial convolution of UNet block. During inference, we use the average of the storyflows computed from the training set as the storyflow $\Delta$.
Model FID $\downarrow$ Char. F1 $\uparrow$ Frm. Acc. $\uparrow$
@@
Table 3: Quantitative comparison with SOTA methods of story visualization on FlintstonesSV.
Model FID $\downarrow$ Char. F1 $\uparrow$ Frm. Acc. $\uparrow$
ContextualStory 13.61 77.24 51.59
--StoryFlow Adapter 14.84 77.09 50.48
--Storyline Contextualizer 15.02 75.42 48.39
    --Spatially-Enhanced Temporal Attention 17.42 71.70 44.83
     --Temporal Convolution 19.69 68.12 39.60
Table 4: Ablation study of the proposed components for the story visualization task on the PororoSV.

### v3

@@ Abstract
Visual storytelling involves generating a sequence of coherent frames from a textual storyline while maintaining consistency in characters and scenes. Existing autoregressive methods, which rely on previous frame-sentence pairs, struggle with high memory usage, slow generation speeds, and limited context integration. To address these issues, we propose ContextualStory, a novel framework designed to generate coherent story frames and extend frames for visual storytelling. ContextualStory utilizes Spatially-Enhanced Temporal Attention to capture spatial and temporal dependencies, handling significant character movements effectively. Additionally, we introduce a Storyline Contextualizer to enrich context in storyline embedding, and a StoryFlow Adapter to measure scene changes between frames for guiding the model. Extensive experiments on PororoSV and FlintstonesSV datasets demonstrate that ContextualStory significantly outperforms existing SOTA methods in both story visualization and continuation. Code is available at https://github.com/sixiaozheng/ContextualStory.
@@ Introduction
Recent text-to-image (T2I) models, such as SD3 (Esser et al. 2024), excel at generating images from text but only produce individual images independently. Although text-to-video (T2V) models like SVD (Blattmann et al. 2023a) and Sora (Brooks et al. 2024) generate coherent videos but often feature simple scene or motion changes. In contrast, this paper focuses on visual storytelling, which comprises generating a sequence of coherent story frames from a textual storyline in story visualization and extending an initial frame from a textual storyline in story continuation. This task has significant potential for educational applications, such as crafting vivid, coherent comics for storybooks. The key challenge is aligning generated frames with sentences while ensuring temporal consistency in characters and scenes. Providing sufficient context is essential due to the limited information in individual sentences.
@@
Figure 1: Story frames generated by our ContextualStory on PororoSV dataset. Red circles highlight character inconsistencies, and blue circles indicate repeated characters. SETA and SC enhance character consistency and scene coherence, achieving superior results compared to AR-LDM.
@@
Many diffusion-based visual storytelling methods use an autoregressive generative approach to capture temporal dependencies based on previous frame-sentence pairs, such as AR-LDM (Pan et al. 2024) and Story-LDM (Rahman et al. 2023). However, these methods face four key limitations: 1) High memory usage due to storing all previous frame-sentence pairs, making longer storyline difficult to handle; 2) Limited context in early frame generation, which may impact frame quality; 3) Slow generation speed due to the sequential nature of the process; 4) Inconsistent frames arise from relying solely on past pairs and neglecting future context, missing the global story context. We address this by exploring how the model can 1) access sufficient frame context and 2) obtain adequate context from the storyline.
@@
To access sufficient frame context, we integrate temporal convolutions and Spatially-Enhanced Temporal Attention (SETA) into the UNet, combining them with the spatial modeling layer. By alternating between spatial and temporal modeling, the model effectively captures spatial dependencies within individual frames and temporal dependencies across frames for comprehensive context. To obtain adequate context from the storyline, we propose the Storyline Contextualizer (SC), which processes the CLIP text embeddings to propagate the context information across sentences, providing sufficient context throughout.
@@
Temporal attention is crucial for visual storytelling as it propagates context across frames along the temporal dimension. However, vanilla temporal attention struggles with significant character movement between frames, as shown in Figure 3(a). To overcome this, we propose SETA that employs a local window mechanism to allow queries to attend to features within local windows of other frames. This effectively captures moving characters and enhances spatial dependency modeling. As shown in Figure 1, SETA improves character consistency and reduces repeated characters, such as Pororo, compared to ContextualStory w/o SETA & SC.
@@
The Storyline Contextualizer enhances the contextual information of storyline embeddings from the CLIP text encoder, which initially contain only sentence-level semantics. By integrating and propagating context across sentences, it generates context-enriched storyline embeddings. The Storyline Contextualizer, a transformer-based network, incorporates self-attention and temporal attention layers to capture both global and temporal dependencies. These enriched embeddings guide the model through a temporally-aligned cross-attention mechanism, ensuring consistent story frames. As shown in Figure 1, ContextualStory improves scene consistency and coherence over ContextualStory w/o SC. Furthermore, compared to AR-LDM, our ContextualStory significantly enhances both characters and scenes consistency.
@@
Additionally, to leverage scene changes between story frames to guide the model, we proposed the StoryFlow Adapter to measure change between story frames. For story continuation, we simply add a convolution layer at the input end of the UNet block to match the size of the first frame latent with the noise latent and then concatenate them.
@@
Our contributions are as follows: (1) ContextualStory Framework: Our novel framework overcomes limitations of existing autoregressive methods, including high memory usage, limited context, slow generation speed, and image inconsistency. (2) Spatially-Enhanced Temporal Attention (SETA): We present the SETA into the UNet model, combining temporal convolutions with spatial modeling to capture both spatial and temporal dependencies, addressing challenges of significant character movement and improving frame consistency. (3) Storyline Contextualizer (SC): A transformer-based network enriches CLIP text embeddings by capturing global and temporal dependencies, ensuring consistent story frames. (4) StoryFlow Adapter: We repurpose this tool to measure scene changes between frames, guiding the model to handle scene transitions more effectively. (5) Extensive experiments on PororoSV and FlintstonesSV datasets demonstrate that our ContextualStory significantly outperforms previous SOTA in visual storytelling.
@@
Visual storytelling. Early methods for story visualization primarily relied on GANs (Goodfellow et al. 2020). StoryGAN (Li et al. 2019) pioneers story visualization using a sequential conditional GAN with a context encoder and dual discriminators to improve narrative and visual coherence. Subsequent works (Song et al. 2020; Li, Torr, and Lukasiewicz 2022; Maharana, Hannan, and Bansal 2021; Maharana and Bansal 2021; Li 2022) improve StoryGAN, while others (Ahn et al. 2023; Chen et al. 2022) adopt Transformer-based methods to enhance character consistency. StoryDALL-E (Maharana, Hannan, and Bansal 2022) extends the story visualization to story continuation with a given initial frame and pre-trained DALL-E (Ramesh et al. 2021). Recently, diffusion models (DM) (Ho, Jain, and Abbeel 2020) have shown success in image generation. Some works (Pan et al. 2024; Rahman et al. 2023; Feng et al. 2023; Song et al. 2024; Liu et al. 2024; Shen and Elhoseiny 2023; Wang et al. 2024) propose an autoregressive diffusion framework based on previous captions and generated frames for consistency. For example, Story-LDM (Rahman et al. 2023) incorporates a visual memory module to capture the context of previous generated images. However, these autoregressive methods are memory-intensive and often fail to capture the global context of the storyline. RCDMs (Shen et al. 2024) is a two-stage model that predicts the embedding of the unknown clip before generating the corresponding images. StoryImager (Tao et al. 2024) is a unified framework for story visualization, continuation, and completion. StoryGPT-V (Shen and Elhoseiny 2023) combines the image generation capability of LDM with the reasoning ability of Large Language Model (LLM) to ensure semantic consistency. TaleCrafter (Gong et al. 2023), Animate-A-Story (He et al. 2023), and AutoStory (Wang et al. 2023c) focus on designing system pipelines for story visualization, all employing LLM to generate storylines. In contrast, our ContextualStory addresses consistency by leveraging SETA to capture complex spatial and temporal dependencies, departing from autoregressive methods.
@@
Text-to-image generation. Recently, significant progress (Rombach et al. 2022; Saharia et al. 2022; Ramesh et al. 2022) has been achieved in T2I generation, primarily due to advancements in DM (Ho, Jain, and Abbeel 2020). Another line of work (Dhariwal and Nichol 2021; Ho and Salimans 2022; Ruiz et al. 2023; Kumari et al. 2023) focuses on flexible and controllable image generation, including ControlNet (Zhang, Rao, and Agrawala 2023), Composer (Huang et al. 2023), IP-Adapter (Ye et al. 2023), and T2I-Adapter (Mou et al. 2024). ControlNet provides a general pipeline for conditioning on both text and image data. The Diffusion Transformer (Peebles and Xie 2023) showcases scalability by replacing UNet with a Transformer, and Pixart-$\alpha$ (Chen et al. 2023) further reduces training costs while achieving superior image quality. However, these methods focus on generating individual images aligned with text and struggle to produce multiple coherent and consistent images in a sequence.
@@
Diffusion models (DM) (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) are generative models that approximate data distributions by iteratively denoising a Gaussian distribution through a reverse process of a Markov Chain. Given a training sample $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$ and add Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to the input in a forward process $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}:={\textstyle\prod_{s=1}^{t}}(1-\beta_{s})$ and $\beta_{1},\ldots,\beta_{T}$ is the variance schedule. The model is trained to approximate the backward process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ by minimizing the mean squared error (MSE) between the predicted and target noise $L_{DM}:=\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|^{2}\right].$
Latent Diffusion Models (LDM) (Rombach et al. 2022) extend DM to high-dimensional data by compressing images into latent space. An encoder $\mathcal{E}$ maps the input $\mathbf{x}$ to a latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})$, , where the forward and backward processes are applied. The denoising network $\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ is trained by minimizing $L_{LDM}:=\mathbb{E}_{\mathcal{E}(\mathbf{x}),\mathbf{c},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})\|^{2}\right],$ where $\mathbf{c}$ denotes conditional signals, such as storyline embeddings. The generated image $\mathbf{\hat{x}}$ is obtained by decoding the denoised latent $\mathbf{z}$ with pre-trained decoder $\mathcal{D}(\mathbf{z})$.
@@
Previous methods based on T2I diffusion models typically use an autoregressive approach, generating story frames sequentially with each frame relying on the preceding frames and captions. However, these methods often fail to capture sufficient storyline context, leading to poor frame consistency. Moreover, the UNet struggles to capture temporal dependencies, while vanilla temporal attention layers are ineffective in addressing significant character movement across frames. To overcome these challenges, as shown in Figure 2, we introduce temporal convolution and SETA into the UNet. These components enable the model to capture contextual information across both spatial and temporal dimensions, allowing it to better handle complex spatial and temporal dependencies. We also propose the Storyline Contextualizer that ensures contextual information propagates to each sentence. Additionally, to address the significant changes in characters and scenes, we introduce the StoryFlow Adapter to quantify these changes and guide the model in generating more coherent visual stories.
Spatially-Enhanced Temporal Attention. In video diffusion models (Zhang et al. 2023), temporal attention layers are often employed to model temporal dependencies. However, unlike video frames with minimal changes and redundant pixels, story frames feature significant character and scene changes. As shown in Figure 3, significant character movement across story frames make it challenging for vanilla temporal attention to capture the same character.
@@
To address this challenge, we propose Spatially-Enhanced Temporal Attention. Assuming the green block within the red-bordered area is the query, the query itself, along with the green blocks covered by the $k\times k$ local window at the same position across other frames (i.e., all the green blocks), form the key and value. Formally, given a hidden state $\mathcal{Z}_{t}=\{\mathbf{z^{1}_{t}},\dots,\mathbf{z^{N}_{t}}\}\in\mathbb{R}^{n\times c\times h\times w}$, where $n=N$ is the number of story frames, and $c$, $h$, $w$ represent the channel, height, and width dimensions of the hidden state, respectively. We first reshape $\mathcal{Z}_{t}$ to $\mathcal{Z}_{t}^{\prime}\in\mathbb{R}^{hw\times n\times c}$, then extract the local window feature $\mathcal{Z}_{t}^{lw^{\prime}}\in\mathbb{R}^{hw\times n^{lw}\times c}$ at each spatial position, where $n^{lw}=(n-1)k^{2}+1$. Subsequently, we compute the query, key, and value and then perform the $\mathrm{Attention}(Q_{T},K_{T},V_{T})$ through Eq. (2).
@@
where $W_{T}^{Q}$, $W_{T}^{K}$, $W_{T}^{V}$ are learnable projection matrices, and $d$ is the feature dimensionality. To ensure a complete local window at the boundaries, we pad $\mathcal{Z}_{t}^{\prime}$ by replicating the boundary features. We utilize rotary positional embedding (RoPE) (Su et al. 2024) as the temporal positional embedding to enable the model understand temporal relationships between frames efficiently.
Storyline Contextualizer. We use the pre-trained CLIP text encoder to independently extract text embeddings for each sentence in the storyline. These embeddings contain the semantic information of the corresponding sentences but lack the global contextual information of the storyline. Directly using these text embeddings to guide the model may result in inconsistent story frames. To address this challenge, we propose the Storyline Contextualizer, which propagates and integrates the contextual information from all text embeddings to generate a context-enriched storyline embedding. As shown in Figure 2, the Storyline Contextualizer is a transformer-based network, each layer contains a self-attention layer, a temporal attention layer, and two Feed-Forward Networks (FFNs).
Given the storyline embedding $\mathcal{C}=\{\mathbf{c^{1}},\dots,\mathbf{c^{N}}\}\in\mathbb{R}^{n\times l\times c_{T}}$ from the CLIP text encoder, where $n=N$ is the number of story sentences, $l$ is the feature sequence length, and $c_{T}$ is the feature dimension. In Storyline Contextualizer, we first reshape the storyline embedding to $1\times nl\times c_{T}$ and apply self-attention, then reshape it to $l\times n\times c_{T}$ for temporal attention. To minimize any adverse effects from additional modules, we zero-initialize the weights of the second FFN in the final layer and incorporate a residual connection, ensuring that the Storyline Contextualizer functions as an identity mapping at the beginning of training. The context-enriched storyline embedding $\mathcal{C}^{\prime}$ guides the model through a cross-attention layer. Unlike T2I and T2V models that compute cross-attention between images/frames and a single text, we adopt a temporally-aligned approach, computing cross-attention between each individual context-enriched text embedding $\mathbf{c^{i}}$ and hidden state $\mathbf{z^{i}_{t}}$ of the corresponding frame.
@@
StoryFlow adapter. To leverage scene changes between story frames to guide the model, we propose the StoryFlow Adapter, inspired by (Jeong, Park, and Ye 2023; Wu et al. 2023; Qing et al. 2023). As shown in Figure 2, the storyflow is computed as the L2 norm of differences between adjacent images, $\delta_{i}=\left\|\mathbf{z_{0}^{i}}-\mathbf{z_{0}^{i+1}}\right\|$, to quantify the image difference. Given $N$ ground truth frames, we calculate the storyflow $\Delta=\{\delta^{1},\dots,\delta^{N-1}\}$. We then encode the storyflow $\Delta$ into a $c$-dimensional embedding using sinusoidal embedding and a zero-initialized FFN. Through linear interpolation, we obtain the storyflow embedding $\Delta^{\prime}\in\mathbb{R}^{N\times c}$. Finally, we add storyflow embedding to timestep embedding and feed them into the spatial convolution of the UNet block. During inference, we use the average of the storyflows computed from the training set as the storyflow $\Delta$.
@@
FlintstonesSV
@@
Table 2: Quantitative comparison with SOTA methods of story visualization on PororoSV and FlintstonesSV.
Solving the Story Continuation Task
For story continuation tasks, besides the storyline, the first frame is provided as an additional input. As shown in Figure 4, we extract the first frame latent $\mathbf{z^{1}}$ and use it as an additional input to all UNet blocks. Within each UNet block, $\mathbf{z^{1}}$ is resized to match the spatial dimensions of the hidden state, then a $1\times 1$ convolution layer adjusts the channels to match the hidden state. Finally, it is concatenated with the hidden state before inputting into the spatial convolution.
Experiment
Experimental Setup
Datasets. We employ two popular benchmark datasets, PororoSV (Li et al. 2019) and FlintstonesSV (Gupta et al. 2018), to evaluate the performance of our model on story visualization and story continuation tasks. PororoSV contains 10,191, 2,334, and 2,208 stories within the train, validation, and test splits, respectively, featuring 9 main characters. FlintstonesSV contains 20,132, 2,071, and 2,309 stories within the train, validation, and test splits, respectively, featuring 7 main characters and 323 backgrounds. Each story in both datasets comprises 5 consecutive story images.
Automatic metrics. To evaluate the quality of generated images, we employ the following three evaluation metrics following previous works (Maharana, Hannan, and Bansal 2022; Pan et al. 2024) in story visualization: (1) Frechet Inception Distance (FID) (Heusel et al. 2017), which measures the distance between feature vectors of ground truth and generated frames; (2) Frame accuracy (Frm. Acc.), which evaluates character matching to ground truth using a fine-tuned Inception-v3 model; (3) Character F1-score (Char. F1), which assesses the quality of generated characters using the same Inception-v3 model as Frm. Acc.
Implementation details. We initialize ContextualStory with the pre-trained Stable Diffusion 2.1-base and fine-tune only the UNet parameters with the AdamW optimizer. Training is performed on 4 NVIDIA A800 GPUs with a batch size of 12, a learning rate of $5\times 10^{-5}$ and 40,000 iterations for PororoSV and 80,000 iterations for FlintstonesSV. The SETA window size is $k=3$, and the SC layer count is 4. During training, we apply classifier-free guidance by randomly dropping input storylines with a 0.1 probability and use the PYoCo mixed noise prior for noise initialization. For inference, we use the DDIM sampler with 50 steps and a guidance scale of 7.5 to generate $256\times 256$ images.